# Learning Encoding-Decoding Direction Pairs to Unveil Concepts of Influence in Deep Vision Networks

### v2

@@ 3.4 Direction Labeling
When learning directions in an unsupervised manner, the name of the concept represented by each encoding-decoding direction pair is unknown. In a real-world setting, the concept name could be identified by manually inspecting the samples that (maximally) activate the decoding direction. However, in a benchmark setting, where interpretability needs to be quantified in terms of metrics and without human involvement, it is necessary to automate the assignment of a concept label to each of the direction pairs. We refer to this process as direction labeling and employ Network Dissection NetworkDissection for its implementation. In essence, Network Dissection assigns a semantic label to each of the concept detectors based on their segmentation performance in a dataset with annotated concepts. Although it was originally proposed as a method to assign labels to each of the basis vectors in the natural latent space basis, Network Dissection is capable of assigning labels to any set of directions after basis change. Despite possible biases against unsupervised learning due to annotation limitations, we adopt and expand on this labeling protocol as a best-effort approach to evaluate the interpretability of our concept detectors and the rest of the unsupervised baselines.
@@ 3.5 Concept Sensitivity Testing
Given the intermediate representation of an image belonging to class $k$ and the direction of concept $c$ in the latent space, RCAV (RCAV) measures the sensitivity of the model to concept $c$ when predicting class $k$. This is accomplished by perturbing the representation towards the direction of the concept with strength $\alpha$, and subsequently comparing the output probability of the network for the same class before and after the perturbation. Subsequently, an overall dataset score in the range [1,1][-1,1] is computed, where zero means inconsistent use of the concept by the model, while extremes indicate consistent positive or negative concept contributions to predicting class $k$. A statistical test compares concept sensitivity against sensitivity towards random directions to ensure significance. In our evaluations, we refer to directions of significant influence in cases where the directions meet the criteria of this statistical significance test. Although the initial work used CAVs to test concept sensitivity, a recent study (PCAV) suggests that the appropriate direction for this purpose is the concept’s encoding direction, an aspect that we consider in our evaluations.
@@ 4 Proposed Method
Figure 4: The proposed method analyzes the latent space to uncover its directional structure. Because many concepts are naturally encoded as specific directions, this process often reveals the encoding-decoding mechanism of meaningful, monosemantic, and highly interpretable concepts. The figure depicts an overview of the method’s components. $\mathcal{L}$ denotes loss terms. Purple indicates contributions of this work, while light gray indicates loss terms from UIBE; CBE.
@@
    Flexible and Sparser Clustering: We lift previous constraints on the concept detectors (such as orthogonality between filters and feature space standardization) in UIBE, enabling more flexible and less redundant (i.e., sparser) clustering. To preserve or even improve interpretability, we incorporate two novel loss terms: Focal Sparsity Loss and Excessively Active Classifier Loss (Section 4.2).
@@ 4.1 Multi-Concept Signal-Distractor Data Model
Figure 5: Left: The learnable parameters of the method $\hat{{\bm{S}}},{\bm{W}},{\bm{b}}$ and intermediate variables $\mathbf{z}_{{\bm{p}}},{\bm{y}}_{{\bm{p}}},{\bm{q}}_{{\bm{p}}}$. Top Right: Feature manipulation and Uncertainty Region Alignment. Bottom Right: Loss terms $\mathcal{L}$ with their dependencies. Purple indicates loss contributions of this work, while light gray indicates loss terms from UIBE; CBE.
@@ 4.2 Losses to Preserve or Improve Interpretability
Excessively Active Classifier Loss ($\mathcal{L}^{eac}$) This loss penalizes excessively large clusters to prevent trivial solutions in which all inputs are assigned to a single cluster. It relies on a hyper-parameter $\rho\in(0,1)$, similar to sparse autoencoders sparse-autoencoder, which sets a proportional bound on cluster size. With ${y}_{{\bm{p}},i}\in[0,1]$ denoting the membership of patch $\bm{p}$} for cluster $i$ (Section 3.3), the unreduced loss formula for the $i$-th cluster is below, with $\gamma>1,\gamma\in{\mathbb{R}}^{+}$ as a sharpening factor, $1-\rho$ normalizing the loss in the range $[0,1]$, and $\bm{p}$} varying across all pixels and image representations in the concept dataset:
@@
Focal Sparsity Loss ($\mathcal{L}^{fs}$) Inspired by Focal Loss (focal-loss), we introduce Focal Sparsity Loss, which places more emphasis on sparsifying the feature-to-cluster assignments of the most challenging patch embeddings. If $\mathcal{L}^{s}_{{\bm{p}}}$ (2) denotes the Sparsity Loss for pixel $\bm{p}$}, the Focal Sparsity Loss is defined as:
@@ 4.3 Signal Vectors as Concept Signal Estimators
In this paragraph, we consider the new data model that was outlined in Section 4.1 and try to validate whether (1) can be used to estimate the signal direction of a concept. Suppose that our objective is to estimate the signal direction of concept $i$ when we have access to a collection of patch embeddings $\{{\bm{x}}_{{\bm{p}}}\}$ and their signal values $\{\alpha_{{\bm{p}},i}\}$. Starting from the approach of PatternNet, it is easy to prove that whenever ${a}_{{\bm{p}},i}$ is independent of all ${a}_{{\bm{p}},j},j\neq i$ and ${\beta}_{{\bm{p}},f}$ the following property holds: $\text{cov}[{\bm{x}}_{{\bm{p}}}-{a}_{{\bm{p}},i}{\bm{s}}_{i},{a}_{{\bm{p}},i}]=0$, which, according to PatternNet directly implies that (1) can be utilized to estimate the direction of the signal.
@@ 4.4 Uncertainty Region Alignment to Discover Meaningful Concepts of Influence
Table 1: Encoding-Decoding Direction Pairs Loss Descriptions. Purple indicates loss contributions of this work, while light gray indicates loss terms from UIBE; CBE
@@ 4.5 Augmented Lagrangian Loss for Effective Direction Learning
The loss terms of the proposed method often conflict with each other. For instance, increasing confidence in the predictions of concept detectors ($\mathcal{L}^{ma}$) conflicts with maximizing the separation margin ($\mathcal{L}^{mm}$), and a large margin may also conflict with sparsity losses ($\mathcal{L}^{s}$ or $\mathcal{L}^{fs}$). Simply using a weighted sum of these losses requires careful architecture and dataset specific tuning of loss weights. The latter can be laborious and only indirectly controls clustering quality, which is mostly determined by the final loss values. To avoid this, we adopt an $\epsilon$-constrained optimization approach and reformulate our objective as a minimization problem with inequality constraints rather than linearly combining all loss terms. In this formulation, hyper-parameter tuning primarily pertains to target loss values, providing direct control over the clustering quality. We use the Augmented Lagrangian formulation (augmented-lagrangian1; augmented-lagrangian2) to solve the following optimization problem:
@@ 5 Encoding-Decoding Direction Pairs in Applications
Moving forward to the application level, in this Section we discuss how to read concept information from the embeddings, how to intervene on the encoding of concepts (Section 5.1) and how to capitalize on the learned direction pairs to obtain local, detailed, and spatially-aware concept explanations (Section 5.2).
@@ 5.1 Reading Concept Information and Intervening on their Encoding
Grouping the first two components in the previous decomposition and exploiting the linearity of GAP, the explanation logit enables the derivation of Concept Contribution Maps (CCMs), which provide a spatially-aware breakdown of the explanation logit in terms of its constituent concept contributions. CCMs differ from previous approaches (IBD; PCA) in that they do not simply localize the concepts in the input image, but they quantify the contributions of each concept across patches to the explanation logit, including both positive and negative contributions.
@@ 6.1 Evaluation of the Signal-Vector Estimator
Based on the synthetic pixel dataset $\{{\bm{x}}_{{\bm{p}}_{1}}\}\cup\{{\bm{x}}_{{\bm{p}}_{2}}\}$ and the ground-truth signal values ${\alpha}_{{\bm{p}},i}$, we put the estimator of PatternNet and our signal vectors under test. The difference between the two estimators is the sub-sampling procedure that we proposed in Section 4.3. Given the ground-truth matrix of signal directions $\bm{S}$} we are able to calculate cosine similarities between the estimated signal directions and the respective ground-truth. The results of this experiment are depicted in Fig. 8 (in blue and green). Since we use the ground-truth signal values, this experiment evaluates the efficacy of signal vectors under perfect input conditions. The figure showcases the effectiveness of our proposed sub-sampling procedure when using (1), as in all cases, signal vectors demonstrate perfect alignment with the ground-truth directions, while in the absence of sub-sampling, the estimation is less reliable. Conclusively, the sub-sampling technique that we proposed in Section 4.3 adapts more precisely to our extended data model.
@@ 7.1 Direction Learning and Evaluation Protocol
We apply our Encoding-Decoding Direction Pairs (EDDP) on the last convolution layer of five different networks: ResNet18 (resnet) trained on Places365 (places365), ResNet50 (resnet) trained on Moments in Time (MiT) (MomentsInTime), as well as EfficientNet (b0) (efficientnet), Inception-v3 (inception), and VGG16 (vgg) trained on ImageNet (imagenet). For all the networks that we study, when learning the direction pairs, we use the Broden (NetworkDissection) concept dataset, except for ResNet50, for which we use Broden-Action BrodenAction. The concept datasets feature dense pixel annotations; Broden includes $1197$ concepts across 63K~63K images in $5$ concept categories (object, part, material, texture, color), while Broden-Action incorporates an additional action category with $210$ labels and 23K~23K more images. We emphasize that when learning the direction pairs, we do not make use of the annotations that complement the concept datasets.
@@
After learning the directions with any unsupervised method, we use the annotations available in the concept datasets to label the directions using Network Dissection (NetworkDissection). As described in Section 3.4, we make a basis change by projecting the network feature activations onto the learned decoding directions (for PCA and EDPP) or by directly utilizing the concept factors learned by NMF. More details on the experimental setup for all methods, including baselines, are provided in Section A.12.
@@ 7.2 Faithfulness Assessment of the Encoding-Decoding Direction Pairs
In this Section, we aim to evaluate the faithfulness of the encoding-decoding direction pairs, i.e. to assess whether the signal value extracted by the decoding direction is indeed encoded in the direction of the signal vector. Although this was validated in synthetic data (Section 6), in this paragraph, we make the same evaluation in the case of real data, in the absence of ground truth. Based on activation maximization (feature-viz-survey), we propose the following novel approach to address this evaluation.
@@
Suppose that we evaluate the faithfulness of a particular direction pair $^{\bm{w}},\hat{{\bm{s}}}$ and let $b$ denote the respective concept detector’s bias. We use the concept detector (${\bm{w}},b$) to collect a set of images containing the concept. Subsequently, with these images as an initialization point, we use Deep Dream (feature-viz-olah) to compute pre-images (pre-image; pre-image2), that is, images that maximize the decoding direction $\bm{w}$} at the spatial locations in which the concept was initially found. In this manner, we start from an image containing the concept and ask Deep Dream to amplify (in the input pixel space) whatever concept the detector can detect. We keep running the activation maximization loop for $K$ iterations and keep a record of how the features (at the locations with the concept) evolve while optimizing. Finally, we use all the recorded features collected during the “dreaming” optimization to fit a parametric line and estimate the direction that best describes feature evolvement while dreaming. We call the direction of this line the dreaming direction (Fig. 9). We conclude by comparing the dreaming direction of the concept for each image with the respective signal vector in terms of cosine similarity.
In Figures 11 and 12, we provide histograms of these cosine similarities when applying our method to ResNet18 and EfficientNet. In all cases, even those without considering $\mathcal{L}^{fso}$, approximately $90\%$ of the signal vectors have cosine similarity with the dreaming directions above 0.7, indicating sufficient faithfulness. The effect of $\mathcal{L}^{fso}$ is to push the distribution to become more Gaussian. To validate the effectiveness of the sub-sampling strategy proposed in Section 4.3, we also provide histograms of cosine similarities between signal vectors and dreaming directions in the absence of sub-sampling. The results are depicted in Fig. 13 where, in that case, the deviation of the signal vectors from the dreaming direction is evident. Example dreaming pre-images for two cases are provided in Fig. 10 and further details and experiments may be found in Sections A.10 and A.11.
@@ 7.4.2 Evaluation of the Decoding Directions: Interpretability
Segmentation We additionally consider the performance of concept detectors in the binary semantic segmentation task and employ two metrics from UIBE. Specifically, let $\phi_{i}(c,\mathcal{K})$ the Intersection Over Union for concept detector $i$ in identifying concept $c$ within the dataset $\mathcal{K}$. Define $c^{\star}_{i}=\text{argmax}_{c}\phi_{i}(c,\mathcal{K}_{t})$, indicating the concept label detected best by concept detector $i$ within the training subset of the dataset ($\mathcal{K}_{t}$). With $\mathcal{K}_{v}$ as the validation subset, we use the following interpretability scores $\mathcal{S}^{1}$ and $\mathcal{S}^{2}$:
@@
Monosemanticity Even though Precision is a natural classification metric that can capture the monosemanticity of the clustering, it has the drawback of relying on the explicit labels that are available in the concept dataset. To overcome this labeling limitation, and inspired by pure-features, we additionally quantify the monosemanticity of the concepts identified by the concept detectors by utilizing the embedding space of CLIP ViT/B-16 (CLIP). In particular, for each concept detector, we pick images from the validation split of the concept dataset for which the detector is most confident about the presence of the concept. We select 100 unique images from the top confident predictions. Subsequently, we crop a rectangle around the area of the concept and obtain the CLIP embedding for this crop. Let $k^{i}_{1}$ and $k^{i}_{2}$ denote image indices in the set of selected images, with $k^{i}_{1},k^{i}_{2}\in\{0,1,...99\}$ and $i$ denoting the $i$-th concept detector. We use the following monosemanticity metric, which measures the average distance between CLIP embeddings (${\bm{e}}_{k^{i}_{1}},{\bm{e}}_{k^{i}_{2}}$) within a cluster, aggregated over all clusters:

### v3

@@ 3.4 Direction Labeling
When learning directions in an unsupervised manner, the name of the concept represented by each encoding-decoding direction pair is unknown. In a real-world setting, the concept name could be identified by manually inspecting the samples that (maximally) activate the decoding direction. However, in a benchmark setting, where interpretability needs to be quantified in terms of metrics and without human involvement, it is necessary to automate the assignment of a concept label to each of the direction pairs. We refer to this process as direction labeling and employ Network Dissection Bau et al. (2017) for its implementation. In essence, Network Dissection assigns a semantic label to each of the concept detectors based on their segmentation performance in a dataset with annotated concepts. Although it was originally proposed as a method to assign labels to each of the basis vectors in the natural latent space basis, Network Dissection is capable of assigning labels to any set of directions after basis change. Despite possible biases against unsupervised learning due to annotation limitations, we adopt and expand on this labeling protocol as a best-effort approach to evaluate the interpretability of our concept detectors and the rest of the unsupervised baselines.
@@ 3.5 Concept Sensitivity Testing
Given the intermediate representation of an image belonging to class $k$ and the direction of concept $c$ in the latent space, RCAV (Pfau et al., 2020) measures the sensitivity of the model to concept $c$ when predicting class $k$. This is accomplished by perturbing the representation towards the direction of the concept with strength $\alpha$, and subsequently comparing the output probability of the network for the same class before and after the perturbation. Subsequently, an overall dataset score in the range [1,1][-1,1] is computed, where zero means inconsistent use of the concept by the model, while extremes indicate consistent positive or negative concept contributions to predicting class $k$. A statistical test compares concept sensitivity against sensitivity towards random directions to ensure significance. In our evaluations, we refer to directions of significant influence in cases where the directions meet the criteria of this statistical significance test. Although the initial work used CAVs to test concept sensitivity, a recent study (Pahde et al., 2024) suggests that the appropriate direction for this purpose is the concept’s encoding direction, an aspect that we consider in our evaluations.
@@ 4 Proposed Method
Figure 4: The proposed method analyzes the latent space to uncover its directional structure. Because many concepts are naturally encoded as specific directions, this process often reveals the encoding-decoding mechanism of meaningful, monosemantic, and highly interpretable concepts. The figure depicts an overview of the method’s components. $\mathcal{L}$ denotes loss terms. Purple indicates contributions of this work, while light gray indicates loss terms from Doumanoglou et al. (2023; 2024).
@@
    Flexible and Sparser Clustering: We lift previous constraints on the concept detectors (such as orthogonality between filters and feature space standardization) in Doumanoglou et al. (2023), enabling more flexible and less redundant (i.e., sparser) clustering. To preserve or even improve interpretability, we incorporate two novel loss terms: Focal Sparsity Loss and Excessively Active Classifier Loss (Section 4.2).
@@ 4.1 Multi-Concept Signal-Distractor Data Model
Figure 5: Left: The learnable parameters of the method $\hat{{\bm{S}}},{\bm{W}},{\bm{b}}$ and intermediate variables $\mathbf{z}_{{\bm{p}}},{\bm{y}}_{{\bm{p}}},{\bm{q}}_{{\bm{p}}}$. Top Right: Feature manipulation and Uncertainty Region Alignment. Bottom Right: Loss terms $\mathcal{L}$ with their dependencies. Purple indicates loss contributions of this work, while light gray indicates loss terms from Doumanoglou et al. (2023; 2024).
@@ 4.2 Losses to Preserve or Improve Interpretability
Excessively Active Classifier Loss ($\mathcal{L}^{eac}$) This loss penalizes excessively large clusters to prevent trivial solutions in which all inputs are assigned to a single cluster. It relies on a hyper-parameter $\rho\in(0,1)$, similar to sparse autoencoders Ng and others (2011), which sets a proportional bound on cluster size. With ${y}_{{\bm{p}},i}\in[0,1]$ denoting the membership of patch $\bm{p}$} for cluster $i$ (Section 3.3), the unreduced loss formula for the $i$-th cluster is below, with $\gamma>1,\gamma\in{\mathbb{R}}^{+}$ as a sharpening factor, $1-\rho$ normalizing the loss in the range $[0,1]$, and $\bm{p}$} varying across all pixels and image representations in the concept dataset:
@@
Focal Sparsity Loss ($\mathcal{L}^{fs}$) Inspired by Focal Loss (Lin et al. (2017)), we introduce Focal Sparsity Loss, which places more emphasis on sparsifying the feature-to-cluster assignments of the most challenging patch embeddings. If $\mathcal{L}^{s}_{{\bm{p}}}$ (2) denotes the Sparsity Loss for pixel $\bm{p}$}, the Focal Sparsity Loss is defined as:
@@ 4.3 Signal Vectors as Concept Signal Estimators
In this paragraph, we consider the new data model that was outlined in Section 4.1 and try to validate whether (1) can be used to estimate the signal direction of a concept. Suppose that our objective is to estimate the signal direction of concept $i$ when we have access to a collection of patch embeddings $\{{\bm{x}}_{{\bm{p}}}\}$ and their signal values $\{\alpha_{{\bm{p}},i}\}$. Starting from the approach of Kindermans et al. (2017), it is easy to prove that whenever ${a}_{{\bm{p}},i}$ is independent of all ${a}_{{\bm{p}},j},j\neq i$ and ${\beta}_{{\bm{p}},f}$ the following property holds: $\text{cov}[{\bm{x}}_{{\bm{p}}}-{a}_{{\bm{p}},i}{\bm{s}}_{i},{a}_{{\bm{p}},i}]=0$, which, according to Kindermans et al. (2017) directly implies that (1) can be utilized to estimate the direction of the signal.
@@ 4.4 Uncertainty Region Alignment to Discover Meaningful Concepts of Influence
Table 1: Encoding-Decoding Direction Pairs Loss Descriptions. Purple indicates loss contributions of this work, while light gray indicates loss terms from Doumanoglou et al. (2023; 2024)
@@ 4.5 Augmented Lagrangian Loss for Effective Direction Learning
The loss terms of the proposed method often conflict with each other. For instance, increasing confidence in the predictions of concept detectors ($\mathcal{L}^{ma}$) conflicts with maximizing the separation margin ($\mathcal{L}^{mm}$), and a large margin may also conflict with sparsity losses ($\mathcal{L}^{s}$ or $\mathcal{L}^{fs}$). Simply using a weighted sum of these losses requires careful architecture and dataset specific tuning of loss weights. The latter can be laborious and only indirectly controls clustering quality, which is mostly determined by the final loss values. To avoid this, we adopt an $\epsilon$-constrained optimization approach and reformulate our objective as a minimization problem with inequality constraints rather than linearly combining all loss terms. In this formulation, hyper-parameter tuning primarily pertains to target loss values, providing direct control over the clustering quality. We use the Augmented Lagrangian formulation (Hestenes, 1969; Bertsekas, 2014) to solve the following optimization problem:
@@ 5 Encoding-Decoding Direction Pairs in Applications
Moving forward to the application level, in this Section, we discuss how to read and intervene on concept information (Section 5.1) and how to capitalize on the learned direction pairs to obtain local, detailed, and spatially-aware concept explanations (Section 5.2).
@@ 5.1 Reading Concept Information and Intervening on their Encoding
Grouping the first two components in the previous decomposition and exploiting the linearity of GAP, the explanation logit enables the derivation of Concept Contribution Maps (CCMs), which provide a spatially-aware breakdown of the explanation logit in terms of its constituent concept contributions. CCMs differ from previous approaches (Zhou et al., 2018; Graziani et al., 2023a) in that they do not simply localize the concepts in the input image, but they quantify the contributions of each concept across patches to the explanation logit, including both positive and negative contributions.
@@ 6.1 Evaluation of the Signal-Vector Estimator
Based on the synthetic pixel dataset $\{{\bm{x}}_{{\bm{p}}_{1}}\}\cup\{{\bm{x}}_{{\bm{p}}_{2}}\}$ and the ground-truth signal values ${\alpha}_{{\bm{p}},i}$, we put the estimator of Kindermans et al. (2017) and our signal vectors under test. The difference between the two estimators is the sub-sampling procedure that we proposed in Section 4.3. Given the ground-truth matrix of signal directions $\bm{S}$} we are able to calculate cosine similarities between the estimated signal directions and the respective ground-truth. The results of this experiment are depicted in Fig. 8 (in blue and green). Since we use the ground-truth signal values, this experiment evaluates the efficacy of signal vectors under perfect input conditions. The figure showcases the effectiveness of our proposed sub-sampling procedure when using (1), as in all cases, signal vectors demonstrate perfect alignment with the ground-truth directions, while in the absence of sub-sampling, the estimation is less reliable. Conclusively, the sub-sampling technique that we proposed in Section 4.3 adapts more precisely to our extended data model.
@@ 7.1 Direction Learning and Evaluation Protocol
We apply our Encoding-Decoding Direction Pairs (EDDP) on the last convolution layer of five different networks: ResNet18 (He et al., 2016) trained on Places365 (Zhou et al., 2017), ResNet50 (He et al., 2016) trained on Moments in Time (MiT) (Monfort et al., 2019), as well as EfficientNet (b0) (Tan and Le, 2019), Inception-v3 (Szegedy et al., 2016), and VGG16 (Simonyan and Zisserman, 2014) trained on ImageNet (Deng et al., 2009). For all the networks that we study, when learning the direction pairs, we use the Broden (Bau et al., 2017) concept dataset, except for ResNet50, for which we use Broden-Action Ramakrishnan et al. (2019). The concept datasets feature dense pixel annotations; Broden includes $1197$ concepts across 63K~63K images in $5$ concept categories (object, part, material, texture, color), while Broden-Action incorporates an additional action category with $210$ labels and 23K~23K more images. We emphasize that when learning the direction pairs, we do not make use of the annotations that complement the concept datasets.
@@
After learning the directions with any unsupervised method, we use Network Dissection (Bau et al., 2017) and the annotations available in the concept datasets to label them. As described in Section 3.4, we make a basis change by projecting the network feature activations onto the learned decoding directions (for PCA and EDPP) or by directly utilizing the concept factors learned by NMF. More details on the experimental setup for all methods, including baselines, are provided in Section A.13, while in Section A.10, we discuss how our evaluation pipeline implicitly captures human judgment.
@@ 7.2 Faithfulness Assessment of the Encoding-Decoding Direction Pairs
In this Section, we aim to evaluate the faithfulness of the encoding-decoding direction pairs, i.e. to assess whether the signal value extracted by the decoding direction is indeed encoded in the direction of the signal vector. Although this was validated in synthetic data (Section 6), in this paragraph, we make the same evaluation in the case of real data, in the absence of ground truth. Based on activation maximization (Nguyen et al., 2019), we propose the following novel approach to address this evaluation.
@@
Suppose that we evaluate the faithfulness of a particular direction pair $^{\bm{w}},\hat{{\bm{s}}}$ and let $b$ denote the respective concept detector’s bias. We use the concept detector (${\bm{w}},b$) to collect a set of images containing the concept. Subsequently, with these images as an initialization point, we use Deep Dream (Olah et al., 2017) to compute pre-images (Mahendran and Vedaldi, 2015; 2016), that is, images that maximize the decoding direction $\bm{w}$} at the spatial locations in which the concept was initially found. In this manner, we start from an image containing the concept and ask Deep Dream to amplify (in the input pixel space) whatever concept the detector can detect. We keep running the activation maximization loop for $K$ iterations and keep a record of how the features (at the locations with the concept) evolve while optimizing. Finally, we use all the recorded features collected during the “dreaming” optimization to fit a parametric line and estimate the direction that best describes feature evolvement while dreaming. We call the direction of this line the dreaming direction (Fig. 9). We conclude by comparing the dreaming direction of the concept for each image with the respective signal vector in terms of cosine similarity.
In Figures 11 and 12, we provide histograms of these cosine similarities when applying our method to ResNet18 and EfficientNet. In all cases, even those without considering $\mathcal{L}^{fso}$, approximately $90\%$ of the signal vectors have cosine similarity with the dreaming directions above 0.7, indicating sufficient faithfulness. The effect of $\mathcal{L}^{fso}$ is to push the distribution to become more Gaussian. To validate the effectiveness of the sub-sampling strategy proposed in Section 4.3, we also provide histograms of cosine similarities between signal vectors and dreaming directions in the absence of sub-sampling. The results are depicted in Fig. 13 where, in that case, the deviation of the signal vectors from the dreaming direction is evident. Example dreaming pre-images for two cases are provided in Fig. 10 and further details and experiments may be found in Sections A.11 and A.12.
@@ 7.4.2 Evaluation of the Decoding Directions: Interpretability
Segmentation We additionally consider the performance of concept detectors in the binary semantic segmentation task and employ two metrics from Doumanoglou et al. (2023). Specifically, let $\phi_{i}(c,\mathcal{K})$ the Intersection Over Union for concept detector $i$ in identifying concept $c$ within the dataset $\mathcal{K}$. Define $c^{\star}_{i}=\text{argmax}_{c}\phi_{i}(c,\mathcal{K}_{t})$, indicating the concept label detected best by concept detector $i$ within the training subset of the dataset ($\mathcal{K}_{t}$). With $\mathcal{K}_{v}$ as the validation subset, we use the following interpretability scores $\mathcal{S}^{1}$ and $\mathcal{S}^{2}$:
@@
Monosemanticity Even though Precision is a natural classification metric that can capture the monosemanticity of the clustering, it has the drawback of relying on the explicit labels that are available in the concept dataset. To overcome this labeling limitation, and inspired by Dreyer et al. (2024b), we additionally quantify the monosemanticity of the concepts identified by the concept detectors by utilizing the embedding space of CLIP ViT/B-16 (Radford et al., 2021). In particular, for each concept detector, we pick images from the validation split of the concept dataset for which the detector is most confident about the presence of the concept. We select 100 unique images from the top confident predictions. Subsequently, we crop a rectangle around the area of the concept and obtain the CLIP embedding for this crop. Let $k^{i}_{1}$ and $k^{i}_{2}$ denote image indices in the set of selected images, with $k^{i}_{1},k^{i}_{2}\in\{0,1,...99\}$ and $i$ denoting the $i$-th concept detector. We use the following monosemanticity metric, which measures the average distance between CLIP embeddings (${\bm{e}}_{k^{i}_{1}},{\bm{e}}_{k^{i}_{2}}$) within a cluster, aggregated over all clusters: